OpenStudy (anonymous):

Mary and Amy had a race, runnning at 6 m/s and 5m/s respectively. How far apart were they after 7 seconds?

OpenStudy (pfenn1):

If the distance Mary ran in 7 seconds is x and the distance Amy ran in 7 seconds is y, then we want to know how far they are apart at 7 seconds or x-y. How would you figure out how far Mary ran in 7 seconds?

OpenStudy (anonymous):

7 x 6= 43m/s 5 x 7= 35m/s

OpenStudy (pfenn1):

x=6 m/s (7 s)=42 m y=5 m/s (7 s)=35 m so the distance they are apart is 7 m. Another way yo calculate is to x-y=6 m/s (7 s)-5 m/s (7s)=(6-5) m/s (7 s)=7 m

OpenStudy (pfenn1):

Close, by the way, 7 x 6 = 42 not 43

OpenStudy (anonymous):

lol aaha

OpenStudy (anonymous):

Thank you :D

OpenStudy (pfenn1):

you're welcome